If a^2+b^2+c^2-2b+6c=-10 then the value of a+b+c=....

Grade:8

2 Answers

Sandipkumar
6 years ago
a^2+b^2+c^2-2b+6c=-10 a^2+b^2+c^2-2b+6c +10=0 a^2+(b^2-2b+1)+(c^2 +6c+ 9)=0a^2 + (b-1)^2 + (c+3)^2 = 0Hence a=0 b=1 & c=-3 or a+b+c=-2
Sandipkumar
6 years ago
a^2+b^2+c^2-2b+6c=-10 a^2+b^2+c^2-2b+6c +10=0 a^2+(b^2-2b+1)+(c^2 +6c+ 9)=0a^2 + (b-1)^2 + (c+3) ^2= 0Hence a=0 b=1 & c=-3 or a+b+c=-2